OpenStudy (anonymous):

I am thinking 70$ Bert is planning to open a savings account that earns 1.6% simple interest yearly. He wants to earn exactly $224 in interest after 2 years. How much money should he deposit? $700 $7,000 $70 $3,584

jimthompson5910 (jim_thompson5910):

Solve for P \[\Large I = P*r*t\] \[\Large 224 = P*0.016*2\]